In a math test, there are 5 questions. the student answered 3 questions. what is the probability that the student can get 3 questions correct?
Answer by addingup(3677)

You have 3 questions that the student answered and each one there is a 50% chance he got it right and a 50% chance he got it wrong:
1/2*1/2*1/2= 1/8 and 1/8= 12.5% chance that he answered all 3 correctly